If (x + 2) is a factor of x^3 − 6x^2 + kx + 10, k =
Answer by stanbon(75887)

If (x + 2) is a factor of x^3 − 6x^2 + kx + 10, k =
-----
If (x+2) is a factor, x = -2 is a zero or root.
-----
Using synthetic division::
-2)....1....-6....k....10
.......1....-4....k-8..|..-2(k-8)+10
------
Solve::
-2(k-8)+10 = 0
-2k + 16 + 10 = 0
-2k = 26 = 0
k = 13
------------
Cheers,
Stan H.
---------------